Smith and Wesson both ran 4 miles. If Smith ran 10% faster than Wesson, how long old it take Smith to run 2.7 miles?

X=Smith=?

Y=Wesson= 24 min
Answer #1

The velocity of Xewwon will be

v_w=\frac{4}{24}=0.1667$ miles/min$

Then velocity of smith will be

v_s=v_s\times\frac{110}{100}=\frac{4}{24}\times\frac{110}{100}=0.1833$ miles/min$

Then time taken by Smith to cover 2.7 miles will be

t=\frac{2.7}{0.1833}=14.73$ min$